OpenStudy (anonymous):

What is S7 for 6 – 24 + 96 – 384 + ... ? I know the answer is 19662..but how do I figure it out?

OpenStudy (whpalmer4):

-24/6 = 96/-24 = -384/96 = see a pattern?

OpenStudy (anonymous):

yeaah!

OpenStudy (anonymous):

thank you :D